# The Extremity Premium: Sentiment Regimes and Adverse Selection in Cryptocurrency Markets

### v1

@@ Abstract
Using the Crypto Fear & Greed Index and Bitcoin daily data, we document that sentiment extremity predicts excess uncertainty beyond realized volatility. Extreme fear and extreme greed regimes exhibit significantly higher spreads than neutral periods—a phenomenon we term the “extremity premium.” Extended validation on the full Fear & Greed history (February 2018–January 2026, N=2,896N=2{,}896) confirms the finding: within-volatility-quintile comparisons show a significant premium ($p<0.001$, Cohen’s $d=0.21$), Granger causality from uncertainty to spreads is strong ($F=211$), and placebo tests reject the null ($p<0.0001$). The effect replicates on Ethereum and across 6 of 7 market cycles. However, the premium is sensitive to functional form: comprehensive regression controls absorb regime effects, while nonparametric stratification preserves them. We interpret this as evidence that sentiment extremity captures volatility-regime interactions not fully represented by parametric controls—consistent with, but not conclusively separable from, the F&G Index’s embedded volatility component. An agent-based model reproduces the pattern qualitatively. The results suggest that intensity, not direction, drives uncertainty-linked liquidity withdrawal in cryptocurrency markets, though identification of “pure” sentiment effects from volatility remains an open challenge.
@@ 1.3 Contributions
    The extremity premium: We document that extreme sentiment regimes (both greed and fear) exhibit elevated uncertainty relative to neutral, even after volatility control. Extreme greed adds +5.5% uncertainty above baseline; extreme fear adds +3.9%. All significant effects survive Bonferroni correction for multiple comparisons.
@@
    Negative result as contribution: Epistemic uncertainty decomposition adds negligible explanatory power ($\Delta R^{2}=0.003$, $p=0.36$), while aleatoric dominates (81.6%). This suggests cryptocurrency sentiment is structurally different from traditional assets—inherently noisy rather than information-asymmetric. The finding redirects research effort from signal refinement to regime detection: simple macro indices suffice.
@@
    SMM-based ABM validation: Rather than informal “stylized facts matching,” we validate the agent-based model via Simulated Method of Moments. The J-test ($p=0.70$) indicates the model is not rejected—it replicates volatility clustering, kurtosis, and spread-volatility correlations without hard-coding them.
@@ 2.1 Agent-Based Market Models
Order book dynamics have received particular attention. Cont et al. (2010) develop a stochastic model matching empirical order flow patterns, while Paddrik et al. (2012) apply ABM to flash crash analysis. Recent work has developed quantitative agent-based models calibrated to real economies, demonstrating that ABMs can match and sometimes outperform mainstream macro models on leverage cycles, bubbles, and crisis dynamics (Farmer, 2025). Complexity-economics applications extend ABM and network models to labor-market transitions and automation shocks, showing how network structure shapes occupational mobility (del Rio-Chanona et al., 2021). A limitation of existing ABM literature for cryptocurrency applications is the treatment of sentiment as either absent or modeled as a single homogeneous signal.
@@ 2.2 Market Microstructure Theory
Avellaneda and Stoikov (2008) extend this framework to high-frequency market making, developing optimal quote-setting strategies that balance inventory risk against adverse selection. Their model shows that market makers should widen spreads when uncertainty about fair value increases—a prediction directly relevant to our uncertainty-driven spread adjustment mechanism. Barucca and Lillo (2017) examine agent reflexivity in price formation, showing how feedback loops between price dynamics and agent beliefs shape microstructure outcomes. Recent financial-computing work characterizes limit order book forecasting and persistence, providing empirical benchmarks for microstructure modeling (Briola et al., 2025a, b). Farzulla (2025c) demonstrates that cryptocurrency liquidity responds differentially to event types: infrastructure failures (exchange collapses, protocol exploits) trigger 65% spread increases versus 11% decreases following regulatory announcements, suggesting that structural uncertainty dominates sentiment-driven uncertainty in determining liquidity provision.
@@ 2.5 Systemic Risk and Macro Signals
Adrian and Brunnermeier (2016) introduce CoVaR, measuring value-at-risk conditional on systemic distress. Gudgeon et al. (2020) analyze the “decentralized financial crisis” of March 2020, documenting how DeFi liquidation cascades amplified market stress. Macro-financial ABMs with explicit banking and interbank markets show how liquidity freezes, policy rules, and network structure shape instability and crisis propagation (Popoyan et al., 2020). Empirical studies of correlation structure and volatility co-movement offer complementary evidence on how shocks propagate across assets, and information filtering networks provide a backbone for extracting sparse dependence structures from high-dimensional markets (Aste, 2025; Samal et al., 2021). Farzulla and Maksakov (2025) develop an aggregated systemic risk index specifically for cryptocurrency markets, combining network topology, liquidity concentration, and cross-exchange contagion measures—providing the macro-level risk signals that complement our micro-level sentiment uncertainty decomposition (live dashboard: asri.dissensus.ai).
@@ 5.4 The Extremity Premium: Extreme Sentiment Amplifies Uncertainty
The “extremity premium”—where extreme sentiment regimes exhibit higher uncertainty than neutral regimes even after controlling for volatility—suggests that market makers face maximum adverse selection risk during sentiment extremes. Pooling extreme greed and extreme fear against neutral yields Cohen’s $d=1.06$ (large effect), indicating the magnitude is economically substantial. When sentiment is directionally intense, informed traders may be exploiting sentiment-driven mispricings, forcing market makers to widen spreads beyond what volatility alone would predict.
@@ 5.9.1 Granger Causality
The reverse direction (spreads $\rightarrow$ uncertainty) is not significant ($F_{3,732}=0.82$, $p=0.49$), supporting the directional interpretation: uncertainty predicts spreads rather than spreads predicting uncertainty.
@@ 5.10.11 Extended Sample Validation: February 2018 – January 2026
[…]
@@ 5.12 ABM Calibration Results
Statistic Real Data Simulation
Daily Return Std 2.49% 1.98%
Return Kurtosis 2.45 11.16
Volatility Clustering (lag-1) 0.30 0.80
Mean Spread 5.0 bps 8.7 bps
The calibrated model successfully reproduces:
@@
    Realistic volatility: Daily return standard deviation within 20% of real data
@@
    Strong volatility clustering: Lag-1 autocorrelation of 0.80, exceeding real data (0.30)
@@
    Reasonable spreads: Mean spread of 8.7 bps
@@ 5.13 Simulated Method of Moments Validation
Beyond informal comparison of stylized facts, we formally validate the ABM using Simulated Method of Moments (SMM), following Grazzini and Richiardi (2015). SMM provides a rigorous test of whether the model is consistent with observed market microstructure.
@@ Results.
Interpretation. The J-test fails to reject the model at any conventional significance level ($p=0.70\gg 0.05$). This provides formal evidence that the ABM is consistent with the data—it replicates key market microstructure moments without having these moments hard-coded into the specification.
Note on J-tests. This SMM J-test ($p=0.70$) evaluates whether the ABM matches market microstructure moments. It differs from the GMM J-test in Section 5.10.8 ($J=75548$, $p<0.001$), which evaluates whether the uncertainty decomposition weights can simultaneously match four distinct uncertainty moments. The SMM passes; the GMM rejects—these are complementary findings about different model components.
@@
6 moments matched, 5 parameters estimated, df = 1. Weighting: identity matrix. $p>0.05$ indicates model is not rejected. The spread-volatility correlation is matched nearly exactly; higher-order autocorrelations show acceptable deviations.
@@ 6.4 Theoretical Implications
The framework extends classic market microstructure models by incorporating sentiment uncertainty decomposition into spread-setting. The key theoretical insight is that sentiment uncertainty—not sentiment level—predicts adverse selection risk.
This finding aligns with emerging evidence that regulatory interventions in cryptocurrency markets affect prices primarily through sentiment channels rather than mechanistic ones. Farzulla (2025c) document that infrastructure events (FTX collapse, Terra/UST) produce 65% spread increases while regulatory events produce 11% spread decreases (p=0.0009p=0.0009)—regulators cannot directly enforce changes to decentralized market structure, so they affect sentiment and expectations rather than infrastructure. If regulatory uncertainty influences spreads through how traders feel about regulation rather than actual structural changes, the epistemic uncertainty component of our decomposition captures a genuine information channel—market participants’ beliefs about regulatory risk, not the risk itself.
@@ 7 Limitations
However, the stratification-based approach—comparing extreme versus neutral spreads within volatility quintiles—yields different conclusions. The pooled within-quintile test remains highly significant ($t=3.36$, $p=0.0008$, Cohen’s $d=0.21$), with the effect concentrated in high-volatility quintiles (Q4: $p=0.017$; Q5: $p=0.055$).
@@
ABM Mechanism Non-Emergence. A legitimate methodological concern is that the agent-based model’s spread-uncertainty correlation is not emergent but architected: market makers explicitly incorporate an uncertainty premium term ($\delta\cdot\sigma_{\text{total}}$) in their quoting logic (Equations 16–17). Finding that simulated spreads correlate with uncertainty therefore confirms implementation fidelity rather than validating the economic mechanism. We acknowledge this limitation directly. The ABM serves a more circumscribed purpose: it provides magnitude calibration and consistency check rather than independent mechanistic validation. Specifically, the SMM procedure validates that when this mechanism operates, the model jointly reproduces four key market microstructure moments—volatility clustering, excess kurtosis, volume persistence, and spread-volatility correlation—without these being hard-coded. The J-test ($p=0.70$) indicates the parametric specification is not rejected, meaning the mechanism’s quantitative magnitude ($\delta=0.18$) is consistent with observed data. What is emergent includes: price trajectories from order flow, volatility clustering from agent feedback loops, fat tails from heterogeneous trader interactions, and regime dynamics. The spread-uncertainty link itself is assumed rather than derived. Setting $\delta=0$ eliminates the correlation entirely (ablation analysis, Section 5.8), indicating the mechanism is load-bearing—but this demonstrates necessity, not sufficiency. The primary evidence for the uncertainty channel remains empirical (Section 5.1); the ABM is an illustrative device that quantifies mechanism magnitude rather than proves mechanism existence.
@@
Causal Identification. While Granger causality tests support a predictive relationship, we do not claim strict causation. ADF tests confirm stationarity, but Granger causality establishes temporal precedence rather than structural causation. Instrumental variables proved weak (first-stage $F<10$), though OLS and IV estimates are nearly identical, suggesting minimal endogeneity bias.
Simulation Limitations. The elevated kurtosis (11.16 vs 2.45) suggests over-coordination in the model. This reflects ABM agent synchronization rather than calibration failure; the SMM targets mean, variance, autocorrelation, and volatility clustering as binding moments, with kurtosis included as a diagnostic rather than a calibration target.
@@
    Uncertainty predicts spreads: Granger causality shows uncertainty predicts spreads ($F=12.79$, $p<0.001$), not vice versa ($F=0.82$, $p=0.49$).

### v2

# Dissensus AI Working Paper Series DAI-2510

@@ Abstract
Using the Crypto Fear & Greed Index and Bitcoin daily data, we document that sentiment extremity predicts excess uncertainty beyond realized volatility. Extreme fear and extreme greed regimes exhibit significantly higher spreads than neutral periods—a phenomenon we term the “extremity premium.” Extended validation on the full Fear & Greed history (February 2018–January 2026, N=2,896N=2{,}896) confirms the finding: within-volatility-quintile comparisons show a significant premium ($p<0.001$, within-quintile stratified Cohen’s $d=0.21$; pooled $d=0.40$), Granger causality from uncertainty to spreads is strong ($F=211$), and placebo tests reject the null ($p<0.0001$). The effect replicates on Ethereum and across 6 of 7 market cycles. However, the premium is sensitive to functional form: comprehensive regression controls absorb regime effects, while nonparametric stratification preserves them. We interpret this as evidence that sentiment extremity captures volatility-regime interactions not fully represented by parametric controls—consistent with, but not conclusively separable from, the F&G Index’s embedded volatility component. An agent-based model reproduces the pattern qualitatively. The results suggest that intensity, not direction, drives uncertainty-linked liquidity withdrawal in cryptocurrency markets, though identification of “pure” sentiment effects from volatility remains an open challenge.
@@ 1.3 Contributions
    The extremity premium: We document that extreme sentiment regimes (both greed and fear) exhibit elevated uncertainty relative to neutral, even after volatility control. Extreme greed adds +5.5 percentage points uncertainty above baseline; extreme fear adds +3.9 percentage points. All significant effects survive Bonferroni correction for multiple comparisons.
@@
    Negative result as contribution: Epistemic uncertainty accounts for only 18.4% of total uncertainty, while aleatoric dominates (81.6%; Table 4). This suggests cryptocurrency sentiment is structurally different from traditional assets—inherently noisy rather than information-asymmetric. The finding redirects research effort from signal refinement to regime detection: simple macro indices suffice.
@@
    SMM-based validation: Rather than informal “stylized facts matching,” we validate a simplified representation of the model’s key mechanisms via Simulated Method of Moments. The J-test ($p=0.70$) indicates the reduced-form model is not rejected—it replicates volatility clustering, kurtosis, and spread-volatility correlations without hard-coding them.
@@ 2.1 Agent-Based Market Models
Order book dynamics have received particular attention. Cont et al. (2010) develop a stochastic model matching empirical order flow patterns, while Paddrik et al. (2012) apply ABM to flash crash analysis. Recent work has developed quantitative agent-based models calibrated to real economies, demonstrating that ABMs can match and sometimes outperform mainstream macro models on leverage cycles, bubbles, and crisis dynamics (Farmer, 2025). Complexity-economics applications extend ABM and network models to labor-market transitions and automation shocks, showing how network structure shapes occupational mobility (del Rio-Chanona et al., 2021). Pangallo and del Rio-Chanona (2024) argue that ABMs require stronger empirical calibration to move beyond stylized-fact matching toward genuine forecasting and policy evaluation. Our calibration to CryptoCompare BTC/USDT data via Simulated Method of Moments (Section 5), with explicit moment-matching targets for volatility clustering, return autocorrelation, and spread dynamics, represents precisely the kind of data-driven grounding they advocate—while our uncertainty decomposition adds a mechanistic layer that purely statistical calibration cannot provide. A limitation of existing ABM literature for cryptocurrency applications is the treatment of sentiment as either absent or modeled as a single homogeneous signal.
@@ 2.2 Market Microstructure Theory
Avellaneda and Stoikov (2008) extend this framework to high-frequency market making, developing optimal quote-setting strategies that balance inventory risk against adverse selection. Their model shows that market makers should widen spreads when uncertainty about fair value increases—a prediction directly relevant to our uncertainty-driven spread adjustment mechanism. Barucca and Lillo (2017) examine agent reflexivity in price formation, showing how feedback loops between price dynamics and agent beliefs shape microstructure outcomes. Recent financial-computing work characterizes limit order book forecasting and persistence, providing empirical benchmarks for microstructure modeling (Briola et al., 2025a, b). A pilot exploration by the author suggests that cryptocurrency liquidity responds differentially to event types, with infrastructure failures (exchange collapses, protocol exploits) producing substantially larger spread increases than regulatory announcements—indicating that structural uncertainty may dominate sentiment-driven uncertainty in determining liquidity provision (Farzulla, 2025e).
@@ 2.5 Systemic Risk and Macro Signals
Adrian and Brunnermeier (2016) introduce CoVaR, measuring value-at-risk conditional on systemic distress. Gudgeon et al. (2020) analyze the “decentralized financial crisis” of March 2020, documenting how DeFi liquidation cascades amplified market stress. Macro-financial ABMs with explicit banking and interbank markets show how liquidity freezes, policy rules, and network structure shape instability and crisis propagation (Popoyan et al., 2020). Empirical studies of correlation structure and volatility co-movement offer complementary evidence on how shocks propagate across assets, and information filtering networks provide a backbone for extracting sparse dependence structures from high-dimensional markets (Aste, 2025; Samal et al., 2021). Farzulla and Maksakov (2025) develop an aggregated systemic risk index specifically for cryptocurrency markets, combining network topology, liquidity concentration, and cross-exchange contagion measures—providing the macro-level risk signals that complement our micro-level sentiment uncertainty decomposition (live dashboard: asri.dissensus.ai). Network-theoretic analysis of financial instability further demonstrates that heterogeneous portfolio allocations can trigger stability-instability phase transitions, with diversification having non-monotonic effects depending on network connectivity (Forer et al., 2025). The bidirectional nature of risk transmission between traditional and decentralized finance—termed ‘crosstagion’ by Aufiero et al. (2025)—further motivates micro-level modeling of how sentiment-driven uncertainty propagates across market structures.
@@ 5.4 The Extremity Premium: Extreme Sentiment Amplifies Uncertainty
The “extremity premium”—where extreme sentiment regimes exhibit higher uncertainty than neutral regimes even after controlling for volatility—suggests that market makers face maximum adverse selection risk during sentiment extremes. Pooling extreme greed and extreme fear against neutral yields Cohen’s $d=1.06$ (large effect), indicating the magnitude is economically substantial. [^1] This pooled $d$ conflates between-regime variance with within-regime variance. The within-quintile stratified effect sizes (Section 7, median $d=0.21$) provide a more conservative measure that controls for volatility-regime heterogeneity. When sentiment is directionally intense, informed traders may be exploiting sentiment-driven mispricings, forcing market makers to widen spreads beyond what volatility alone would predict.
@@ 5.9.1  Granger Causality
The reverse direction (spreads $\rightarrow$ uncertainty) is not significant ($F_{3,732}=0.82$, $p=0.49$), supporting the directional interpretation: uncertainty predicts spreads rather than spreads predicting uncertainty. However, the extended sample (Section 5.10.11) reveals significant bidirectional causality at lags 1–4, suggesting a weaker reverse channel that activates during structural breaks; the forward direction remains dominant by two orders of magnitude in F-statistic.
@@ 5.10.11  Extended Sample Validation: February 2018 – January 2026
Interpretation. The extended sample provides overwhelming evidence for the extremity premium. With $p=2.7\times 10^{-14}$ and a 95% CI that excludes zero by a wide margin, the null hypothesis that extreme and neutral regimes have equal spreads can be definitively rejected. The effect replicates across bull markets (2020, 2021, 2024–25), bear markets (2018, 2022), and on both BTC and ETH. Granger causality is nearly seven-fold stronger with the larger sample (F = 211 vs. 31). Both placebo tests now achieve $p<0.0001$.
@@
Bidirectional Granger Causality in the Extended Sample. While the main sample (739 days) shows clean unidirectional Granger causality—uncertainty predicts spreads but not vice versa (all reverse p>0.48p>0.48)—the extended sample reveals a more nuanced picture. The reverse direction (spreads $\rightarrow$ uncertainty) becomes significant at lags 1–4 ($p=0.040$, $0.005$, $0.003$, $0.012$ respectively), though not at lag 5 ($p=0.089$). The forward direction (uncertainty $\rightarrow$ spreads) remains overwhelmingly significant at all lags ($F>50$, $p<10^{-46}$). The asymmetry in magnitude persists: the forward F-statistics (51–211) dwarf the reverse (1.9–5.4), indicating the predominant predictive direction is unchanged. We interpret the bidirectional finding as reflecting the extended sample’s coverage of major structural breaks—the March 2020 COVID crash, the May 2021 China mining ban, and the November 2022 FTX collapse—during which extreme spread dislocations plausibly fed back into uncertainty measures. In shorter, more homogeneous samples, this feedback channel is absent. This finding qualifies but does not overturn the directional interpretation: uncertainty is the primary driver, with a weaker reverse channel that activates during market crises.
@@ 5.12 ABM Calibration Results
Statistic Real Data Initial Calibration Recalibrated
Daily Return Std 2.49% 1.98% 1.14%
Return Kurtosis 2.45 11.16 4.49
Volatility Clustering (lag-1) 0.30 0.80 0.05
Mean Spread 5.0 bps 8.7 bps 4.3 bps
Table 39: *
The initial calibration used 3 market makers with wide base spreads (15 bps); the recalibrated version uses 5 market makers with competitive spreads (7 bps) and higher noise trader activity, yielding more realistic spread magnitudes and kurtosis. Volatility clustering decreases because the Mesa ABM lacks explicit GARCH dynamics—the initial calibration’s 0.80 reflected regime persistence in a small sample (30 daily returns) rather than true volatility clustering. The recalibrated 0.05 uses session-level aggregation (150+ observations per run) and is below the empirical target, reflecting the ABM’s simplified volatility mechanism.
The recalibrated model improves on several dimensions:
@@
    Realistic spreads: Mean spread of 4.3 bps (within the 2–5 bps range for major exchanges), improved from 8.7 bps in the initial calibration
@@
    Improved kurtosis: Return kurtosis of 4.49, within the typical range for cryptocurrency returns (4–8), reduced from 11.16
@@
    Volatility clustering: Lag-1 autocorrelation of 0.05, below the empirical target of 0.30. The Mesa ABM does not implement GARCH-type variance persistence; observed clustering emerges solely from sentiment regime transitions. This is a known limitation of the model specification
@@ 5.13 Simulated Method of Moments Validation
Beyond informal comparison of stylized facts, we formally validate the model using Simulated Method of Moments (SMM), following Grazzini and Richiardi (2015). SMM provides a rigorous test of whether the model is consistent with observed market microstructure. For computational tractability, the SMM estimation uses a simplified representation of the key mechanisms—a reduced-form chartist-fundamentalist model with GARCH-like volatility and uncertainty-dependent spreads—rather than the full Mesa ABM described in Section 3.4. This simplified model captures the same economic channels (uncertainty-sensitive spread widening, heterogeneous trader behavior, regime-dependent dynamics) but with tractable parameter estimation. The Mesa ABM’s computational cost per evaluation (requiring order-book simulation with discrete agents) makes direct SMM infeasible within reasonable time constraints. To be explicit: the SMM validates the reduced-form economic mechanism—that uncertainty drives spread widening at magnitudes consistent with data—not the full agent specification with 200+ heterogeneous agents, limit-order-book dynamics, and emergent price formation. The J-test ($p=0.70$) indicates that the parametric uncertainty channel is not rejected by the data; it does not validate the Mesa ABM’s richer agent ecology.
@@ Results.
Interpretation. The J-test fails to reject the model at any conventional significance level ($p=0.70\gg 0.05$). This provides formal evidence that the simplified model’s mechanisms are consistent with observed market microstructure—it replicates key moments without having them hard-coded into the specification. The result supports the economic plausibility of uncertainty-sensitive spread widening but should be interpreted as validating the reduced-form mechanism rather than the full agent-based specification.
Note on J-tests. This SMM J-test ($p=0.70$) evaluates whether the simplified model matches market microstructure moments. It differs from the GMM J-test in Section 5.10.8 ($J=75548$, $p<0.001$), which evaluates whether the uncertainty decomposition weights can simultaneously match four distinct uncertainty moments. The SMM passes; the GMM rejects—these are complementary findings about different model components.
@@
6 moments matched, 5 parameters estimated, df = 1. Weighting: identity matrix. $p>0.05$ indicates model is not rejected. The spread-volatility correlation is matched nearly exactly; higher-order autocorrelations show acceptable deviations. The empirical kurtosis target (2.23) differs from Table 12 (2.45) because the reduced-form SMM model uses a slightly different sample window for moment computation; both values indicate moderate leptokurtosis consistent with cryptocurrency return distributions.
@@ 6.4 Theoretical Implications
The framework extends classic market microstructure models by incorporating sentiment uncertainty decomposition into spread-setting, contributing to the broader complexity economics research programme that deploys agent-based, network, and dynamical systems methods to address challenges where equilibrium approaches fall short (Bednar et al., 2025). The key theoretical insight is that sentiment uncertainty—not sentiment level—predicts adverse selection risk.
This finding aligns with emerging evidence that regulatory interventions in cryptocurrency markets affect prices primarily through sentiment channels rather than mechanistic ones. A pilot exploration by the author finds that infrastructure events (FTX collapse, Terra/UST) produce substantially larger spread increases than regulatory events, which may in fact decrease spreads—suggesting that regulators cannot directly enforce changes to decentralized market structure, affecting sentiment and expectations rather than infrastructure (Farzulla, 2025e). If regulatory uncertainty influences spreads through how traders feel about regulation rather than actual structural changes, the epistemic uncertainty component of our decomposition captures a genuine information channel—market participants’ beliefs about regulatory risk, not the risk itself.
@@ 7 Limitations
However, the stratification-based approach—comparing extreme versus neutral spreads within volatility quintiles—yields different conclusions. The pooled within-quintile test remains highly significant ($t=3.36$, $p=0.0008$; the aggregate pooled comparison yields $d=0.40$), with the effect concentrated in high-volatility quintiles (Q4: $p=0.017$; Q5: $p=0.055$). We report a within-quintile stratified Cohen’s $d=0.21$, computed as the median of the five quintile-level effect sizes ($d_{q}=0.16,0.21,0.21,0.16,0.23$); the $n$-weighted mean is $0.20$. We prefer the median as a robust summary that is not dominated by the largest quintile cells.
@@
ABM Mechanism Non-Emergence. A legitimate methodological concern is that the agent-based model’s spread-uncertainty correlation is not emergent but architected: market makers explicitly incorporate an uncertainty premium term ($\delta\cdot\sigma_{\text{total}}$) in their quoting logic (Equations 16–17). Finding that simulated spreads correlate with uncertainty therefore confirms implementation fidelity rather than validating the economic mechanism. We acknowledge this limitation directly. The ABM serves a more circumscribed purpose: it provides magnitude calibration and consistency check rather than independent mechanistic validation. Specifically, the SMM procedure validates that when this mechanism operates, a simplified representation of the model jointly reproduces four key market microstructure moments—volatility clustering, excess kurtosis, volume persistence, and spread-volatility correlation—without these being hard-coded (see Section 5.13 for the distinction between the Mesa ABM and the SMM estimation model). The J-test ($p=0.70$) indicates the parametric specification is not rejected, meaning the mechanism’s quantitative magnitude ($\delta=0.18$) is consistent with observed data. What is emergent includes: price trajectories from order flow, volatility clustering from agent feedback loops, fat tails from heterogeneous trader interactions, and regime dynamics. The spread-uncertainty link itself is assumed rather than derived. Setting $\delta=0$ eliminates the correlation entirely (ablation analysis, Section 5.8), indicating the mechanism is load-bearing—but this demonstrates necessity, not sufficiency. The primary evidence for the uncertainty channel remains empirical (Section 5.1); the ABM is an illustrative device that quantifies mechanism magnitude rather than proves mechanism existence.
@@
Causal Identification. While Granger causality tests support a predictive relationship, we do not claim strict causation. ADF tests confirm stationarity, but Granger causality establishes temporal precedence rather than structural causation. Instrumental variables proved weak (first-stage $F<10$), though OLS and IV estimates are nearly identical, suggesting minimal endogeneity bias. Importantly, the extended sample reveals bidirectional Granger causality at lags 1–4 (Section 5.10.11), suggesting a weak reverse channel (spreads $\rightarrow$ uncertainty) that emerges during structural breaks. The directional claim is strongest in the main sample; the extended sample supports a predominantly forward relationship with crisis-period feedback.
Simulation Limitations. The initial Mesa ABM calibration exhibited elevated kurtosis (11.16 vs 2.45) and volatility clustering (0.80 vs 0.30 empirical), reflecting agent synchronization and small-sample measurement artifacts respectively. Recalibration with increased market maker competition and higher trading activity reduced kurtosis to 4.49 and spreads to 4.3 bps, but volatility clustering fell to 0.05—below the empirical target. This reflects the ABM’s lack of explicit GARCH dynamics; volatility persistence emerges from sentiment regimes rather than autoregressive variance. Table 39 reports both calibration runs for transparency.
@@
    Uncertainty predicts spreads: Granger causality shows uncertainty predicts spreads ($F=12.79$, $p<0.001$); the reverse direction is not significant in the main sample ($F=0.82$, $p=0.49$) but becomes significant at lags 1–4 in the extended sample, consistent with crisis-period feedback (see Section 5.10.11).